Apr 15

6:20 Starting up LUCI on the remote control room at Steward.

7:26 Start collimating the telescope.

7:35 Collimation finished. Seeing=0.85"

7:35 Going to telluric star HIP53735. Problem with the 1st script; Switching to another.d

8:06 Start integrating for telluric star HIP53735 J-band/H-band MOS for A851.

8:19 Moving to A851.

8:54 Mirror panicked.

9:10 Mirror recovered. Collimation started. Reacquiring A851. Total downtime due to mirror panic = 33min.

9:28 Started integration on A851 (H-band spectra). Seeing ~0.5"-0.6".

11:48 Handing the control over to Benjamin at LBT.

12:00 Moving to the next target CLJ1226. Seeing~0.8".

12:21 Starting integration in the H-band with CLJ1226.

1:25 The probe failed to move, offset correctly, lost guiding (frame 93)

1:30 Geno checked guiding from the guide probe failure

1:35 CLJ1226 K arc acquisition

1:49 CLJ1226 K arc start integration

2:37 R1532 acquisition

3:22 R1532 start integration, mask alignment was difficult, offsets computed incorrectly by LUCI software

3:37 Accidentally launched acquisition script, recovered, now integrating, seeing ~0.8"

5:12 done with science on R1532, moving to telluric HIP092548

5:28 done with telluric HIP092548 using R1532 mask.

5:32 can't take telluric with long-slit because sun is rising soon and we are facing East.

6:07 taking flats

6:55 done with calibs

-- %USERSIG{Eiichi Egami - 2014-04-15}%

Comments


 
Topic revision: r5 - 16 Apr 2014, GregoryLeeWalth
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback